5. Define a Markov Chain on S-1,2,3,..) with transition probabilities Pi i+1 (a) Is the MC irreducible? (b) Are the states positive recurrent? (c) Find the invariant distribution
Answer #1

(a)

The probability distribution matrix can be written as,

1/2 1/2 00 2/3 1/3 i/i+1 0 0.. 1/i+1

We see that we can reach to state 1 from each state and from state 1, we can reach to all the states. Thus, all states can communicate to each other and the MC is irreducible.

(b)

A positive recurrent Markov chain is one where the average number of steps needed to return to a state is a finite number.

For state infty, in the next transaction it can either go to state 1 or state 00 +1. From state 00 +1, it can reach to state infty only through state 1. Now from state 1, average number of steps needed to reach state infty is not finite (as there are infinite number of states in between). Thus, the states are not positive recurrent.

(c)

Let pi = [a_1, a_2, ...., a_i, a_{i+1},....] be the invariant distribution vector.

Then, pi P = pi

and a a2+.... + ai ai+1 +...-1 ----(1)

and

a1/2 + 2a2/3+.... +ia,/(i +1) +..-a

aDRightarrow a_2 = a_1/2!

a_2/3 = a_3   3a1/2%3

.....

di-1   Rightarrow a_{i} = a_1/i!

a_i/(i+1) = a_{i+1}Rightarrow a_{i+1} = a_1/(i+1)!

---

From (1), and Substituting values of a2, a3, ..., ai,... we get

a a2+.... + ai ai+1 +...-1

a_1+ a_1/2! + ....+ a_i /i!+ a_{i+1}/(i+1)! + ... = 1

a_1(1+ 1/2! + ....+ 1 /i!+ 1/(i+1)! + ...) = 1

11 (e-1)=1 where e is the exponential(1)

a_1= 1/(e - 1)

Hence,

Rightarrow a_2 = 1/2!(e-1)

..

Rightarrow a_{i} = a_1/i!(e-1)

Thus, the invariant distribution is,

pi = [1/1!(e - 1), 1/2!(e - 1) ...., 1/i!(e - 1),....]